OpenStudy (anonymous):

How do I simplify ((x^(a+b))^(a-b))/((x^(a-2b))^(a+2b))

OpenStudy (lilai3):

congrats on you first question @OdinMW

OpenStudy (anonymous):

Thanks, Its a fraction. Im this is the only way I could write it out. Sorry for all the brakets.

OpenStudy (lilai3):

ikr looks hard @ 1st glance

OpenStudy (campbell_st):

well you need to use the power of a power for the numerator and denominator \[(x^a)^b = x^{a \times b}\] once you have used the rule.. look for common factors... to cancel

OpenStudy (anonymous):

Would the first part be x^(a^2-b^2)?

OpenStudy (campbell_st):

thats correct, well done

OpenStudy (anonymous):

If i took (x^(a+b))^(a-b)

OpenStudy (campbell_st):

yep the index law for power of a power says multiply the powers.

OpenStudy (anonymous):

Is the correct answer 4?

OpenStudy (lilai3):

i think that's what i got....it took me a really long time tho -_- want me to explain...nvm cuz i saw campbell has already done so :)

OpenStudy (anonymous):

I hope its right :) thanks for confirming!

OpenStudy (campbell_st):

thats impossible... this is what you have so far \[\frac{x^{(a^2 - b^2)}}{(x^{(a - 2b})^{(a + 2b)}}\] you need to use the power of a power on the numerator...

OpenStudy (anonymous):

So on the bottom I have \[x ^{a ^{2}-4b ^{2}}\]

OpenStudy (anonymous):

Which makes it \[\frac{ x ^{a ^{2}-b ^{2}} }{ x ^{a ^{2}-4b ^{2}} }\]

OpenStudy (campbell_st):

thats correct

OpenStudy (campbell_st):

so now you have your simplified fraction...

OpenStudy (anonymous):

.... wait

OpenStudy (campbell_st):

now you need to use the power rule for dividing \[\frac{x^a}{x^b} = x^{a - b}\]

OpenStudy (campbell_st):

I have to go... but you need to simplify \[x^{(a^2 - b^2)-(a^2 - 4b^2)}\] take care with the subtraction...

OpenStudy (anonymous):

Would it than be \[x ^{a ^{2}-b ^{2}-a ^{2}-({4b) ^{2}}}\] So \[x ^{-b ^{2}+4b ^{2}}\] So \[x ^{3b ^{2}}\] ?

OpenStudy (campbell_st):

thats correct... well done

OpenStudy (anonymous):

Thanks!!!!!!

OpenStudy (campbell_st):

glad to help